# Cost-Effective Text Clustering with Large Language Models

### v1

@@ 7.2. Text Clustering Performance
Table 3 reports the ACC and NMI scores achieved by all the competitors and our TECL on the five datasets. The best and second-best results are highlighted in blue, and darker shades indicate better clustering. The best baselines are underlined. Specifically, we can make the following observations. First, TECL consistently and considerably outperforms all the baselines over all datasets, even including the semi-supervised method PO-PCKMeans. For example, on Tweet and CLINC datasets, TECL can achieve a significant improvement of $7.29\%$ and $5.56\%$ in terms of ACC compared to the best baseline results, respectively. The remarkable gains validate the effectiveness of our proposed framework, which leverages greedy edge/triangle selection algorithms and LLMs to create informative pair constraints, followed by weighted constrained clustering. Amid the four versions of TECL, TriangleLLM is always superior to EdgeLLM since the former yields more pair constraints compared to the latter under the same query budget $Q$, as pinpointed in Section 4. On BBC News, Tweet, and Bank77 datasets, it can be seen that WCSC produces better clusters than WCKMeans, while the latter outperforms the former on Reddit and CLINC. Unless otherwise specified, we refer to TECL with TriangleLLM, WCSC as TECL for simplicity in the remaining experiments.
Another key observation we can make from Table 3 is that the performance of most embedding-based approaches is subpar, except Instructor-Large and OpenAI-GPT, and Instructor-Large is even superior or comparable to OpenAI-GPT on multiple datasets, which reveals that small PLMs can also output high-quality text embeddings for clustering. However, both of them are still inferior to other LLM-based approaches, including ClusterLLM, PO-PCKMeans, and our TECL. The results manifest that the direct adoption of the embeddings from LLMs for text clustering is less than satisfactory, whereas using the explicit correlations between text instances from LLMs shows higher effectiveness, which is consistent with the observation in other studies (Silwal et al., 2023; Zhang et al., 2023a).
Table 5. TECL with various selection strategies.
@@ 7.3. Performance when Varying Query Budgets
In this set of experiments, we further study the text clustering performance of TECL and other LLM-based baselines ClusterLLM and PO-PCKMeans when the query budget $Q$ is varied. Fig. 4 depicts their ACC results on BBC News and Tweet when varying $Q$ in $\{\frac{\Omega}{10},\frac{\Omega}{5},\frac{\Omega}{2},\Omega,2\Omega\}$. It can be observed that ClusterLLM is unable to leverage the feedback from LLMs effectively, as its performance slightly decreases when $Q$ increases. A similar issue can be observed in PO-PCKMeans, whose ACC scores undergo a remarkable decline when $Q$ is increased from $\Omega$ (resp. $\frac{\Omega}{5}$) to $2\Omega$ (resp. $\frac{\Omega}{2}$) on BBC News (resp. Tweet). In sum, both baselines fail to obtain notable improvements when expanding the budget for querying LLMs.
@@
Figure 5. TECL with various constrained clustering methods
By contrast, a consistent and conspicuous performance rise can be observed for TECL on both tested datasets, and the ACC scores are always considerably higher than those by ClusterLLM and PO-PCKMeans under the same query budget. In particular, when the query budget $Q=\frac{\Omega}{10}$, TECL achieves the clustering performance that is even superior to those attained by ClusterLLM and PO-PCKMeans with $Q=2\Omega$ on both datasets, indicating the high query effectiveness of TECL.
@@
Method Weight BBC News Tweet
ACC NMI Acc NMI
WCSC None 90.43 73.90 65.33 87.84
ESS 72.40 65.31 66.55 87.22
SESS 71.37 63.35 66.34 87.17
LESS 71.15 64.36 65.33 86.85
SEC 69.48 63.82 67.27 87.63
SSEC 70.07 65.21 67.72 87.28
IPMI 73.12 65.91 68.89 88.01
PMI (Ours) 91.01 75.22 73.10 89.08
WCKmeans None 89.80 72.38 65.66 87.84
ESS 90.11 73.41 64.00 87.52
SESS 89.98 73.08 63.63 87.41
LESS 90.11 73.33 64.56 87.00
SEC 88.85 70.83 63.71 86.90
SSEC 89.03 70.93 63.63 87.04
IPMI 90.47 74.03 66.63 87.45
PMI (Ours) 90.65 74.25 72.25 88.60
@@ 7.4. Ablation Study and Component Analysis
This section presents ablation studies and the efficacy of each ingredient in TECL.
Ablation Study. We conduct an ablation study to explore the effects of WCSC proposed in Section  4 and TriangleLLM in Section  6.
As shown in Table 4, the removal of WCSC results in a decrease in performance, especially on Tweet (ACC falls from 73.10% to 65.33%). This indicates that the weighting of constraints helps in managing noisy or less informative pairs. Additionally, the absence of TriangleLLM also causes a drop in performance, especially on BBC News (ACC declines from 91.01% to 88.72%), highlighting the importance of selecting high-quality constraints. These findings suggest that both components play a significant role.
Edge/Triangle Selection Strategies. We study three strategies in TriangleLLM to select triangles for querying LLMs, where the Max (resp. Min) is to maximize (resp. minimize) the SEC scores, whilst the Random generates triangles randomly. Table 5 shows the Max strategy as implemented by Algorithm 4 achieves the best performance (e.g., 91.01% ACC on BBC News). In comparison, the Min and Random strategies show significantly inferior performance due to selecting low-SEC or random triangles. The results validate the effectiveness of our greedy selection scheme proposed in Section 6.
Constrained Clustering Methods. Fig. 5 presents the clustering performance of TECL with various constrained clustering methods, including WCSC (ours), CSC (Wang et al., 2014), PCKMeans (Basu et al., 2004), COPKMeans (Bibi et al., 2023), CCPCKMeans (Rujeerapaiboon et al., 2019), and MPCKMeans (Bilenko et al., 2004) on BBC News and Tweet. Our WCSC that assign personalized weights to constraints achieves the highest accuracy (91.01% on BBC and 73.10% on Tweet), indicating its superiority to exploit pairwise constraints.
Weighting Schemes for Constraints. In Table 6, we report the results of TECL when using various weighting schemes in our WCSC and WCKmeans on BBC News and Tweet. The additional weighting schemes for comparison are listed in the following table:
@@
Table 6 shows that improper weights for constraints lead to a significant adverse impact on the performance in both WCSC and WCKmeans. Amid all weighting schemes, the PMI defined in Section 4 consistently achieves the best results. For example, on Tweet, using PMI in WCSC improves over its unweighted version (i.e., None) by a large margin of $7.77\%$, while the best competitor IPMI can only attain an improvement of $3.56\%$.
Table 7. Clustering performance with various text encoders.
Encoder Method BBC News Tweet
ACC NMI ACC NMI
Sentence-BERT SCCL 82.00 58.33 53.60 72.80
ClusterLLM 80.08 54.13 57.47 83.74
PO-PCKMeans 82.70 57.43 61.39 86.12
TECL 84.45 61.65 78.48 89.30
Instructor-Large SCCL 83.60 62.00 36.50 68.50
ClusterLLM 88.40 69.73 65.81 88.61
PO-PCKMeans 89.53 72.16 65.58 88.20
TECL 91.01 75.22 73.10 89.08
Various Text Encoders. Table 7 conveys the clustering performance of TECL and recent baselines with two strong text encoders. The results show the superior effectiveness of TECL: (1) TECL achieves state-of-the-art performance in various encoders, achieving 91. 01% ACC on BBC News with Instructor-Large and 78.48% ACC on Tweet with Sentence-BERT, surpassing the second-best baselines. (2) Under the same encoder, TECL consistently outperforms competitors: with Sentence-BERT, TECL exceeds the second-best method (PO-PCKMeans) with a gain of 1.75% in ACC on BBC News; with Instructor-Large, TECL surpasses ClusterLLM by an improvement of 7.29% in ACC on Tweet. These gains, coupled with superior NMI scores, validate TECL’s ability to leverage pretrained embeddings and cherry-pick constraints without fine-tuning.
GPT-4oGPT-4o-Mini

### v2

@@ 7.2. Text Clustering Performance
Table 3 reports the ACC and NMI scores achieved by all the competitors and our Cequel on the six datasets. The best results are bolded, and the best baselines are underlined. Specifically, we can make the following observations. First, Cequel consistently and considerably outperforms all the baselines over all datasets, even including the semi-supervised method PO-PCKMeans. For example, on Tweet and CLINC datasets, Cequel can achieve a significant improvement of $7.29\%$ and $5.56\%$ in terms of ACC compared to the best baseline results, respectively. The remarkable gains validate the effectiveness of our proposed framework, which leverages greedy edge/triangle selection algorithms and LLMs to create informative pair constraints, followed by weighted constrained clustering. Amid the four versions of Cequel, TriangleLLM is always superior to EdgeLLM since the former yields more pair constraints compared to the latter under the same query budget $Q$, as pinpointed in § 4. On BBC News, Tweet, and Bank77 datasets, it can be seen that WCSC produces better clusters than WCKMeans, while the latter outperforms the former on Reddit, CLINC, and Massive. Unless otherwise specified, we refer to Cequel with TriangleLLM, WCSC as Cequel for simplicity in the remaining experiments.
Another key observation we can make from Table 3 is that the performance of most embedding-based approaches is subpar, except Instructor-Large and OpenAI-GPT, and Instructor-Large is even superior or comparable to OpenAI-GPT on multiple datasets, which reveals that small PLMs can also output high-quality text embeddings for clustering. However, both of them are still inferior to other LLM-based approaches, including ClusterLLM, PO-PCKMeans, and our Cequel. The results manifest that the direct adoption of the embeddings from LLMs for text clustering is less than satisfactory, whereas using the explicit correlations between text instances from LLMs shows higher effectiveness, which is consistent with the observation in other studies (Silwal et al., 2023; Zhang et al., 2023a).
Table 5. Cequel with various selection strategies.
@@
Bank77 Max (Ours) 69.03 83.20
Min 65.84 82.58
Random 66.49 83.10
Reddit Max (Ours) 57.78 64.25
Min 57.36 64.03
Random 57.22 63.83
@@ 7.3. Performance when Varying Query Budgets
In this set of experiments, we further study the text clustering performance of Cequel and other LLM-based baselines ClusterLLM and PO-PCKMeans when the query budget $Q$ is varied. Fig. 4 depicts their ACC results on BBC News, Tweet, Bank77, and Reddit when varying $Q$ in $\{\frac{\Omega}{10},\frac{\Omega}{5},\frac{\Omega}{2},\Omega,2\Omega\}$. It can be observed that ClusterLLM is unable to leverage the feedback from LLMs effectively, as its performance slightly decreases when $Q$ increases on BBC News. A similar issue can be observed in PO-PCKMeans, whose ACC scores undergo a remarkable decline when $Q$ is increased from $\Omega$ (resp. $\frac{\Omega}{5}$) to $2\Omega$ (resp. $\frac{\Omega}{2}$) on BBC News (resp. Tweet). In sum, both baselines fail to obtain notable improvements when expanding the budget for querying LLMs.
@@
0.570.600.630.660.690.72ACC
((c)) Bank77
0.500.520.540.560.580.60ACC
((d)) Reddit
Figure 5. Varying constrained clustering methods in Cequel.
By contrast, a consistent and conspicuous performance rise can be observed for Cequel on all tested datasets, and the ACC scores are always considerably higher than those by ClusterLLM and PO-PCKMeans under the same query budget. In particular, when the query budget $Q=\frac{\Omega}{10}$, Cequel achieves the clustering performance that is even superior to those attained by ClusterLLM and PO-PCKMeans with $Q=2\Omega$ on BBC News and Tweet, indicating the high query effectiveness of Cequel.
@@
Weight BBC News Tweet Bank77 Reddit
ACC NMI ACC NMI ACC NMI ACC NMI
WCSC None 90.43 73.90 65.33 87.84 67.12 82.65 56.30 63.74
ESS 72.40 65.31 66.55 87.22 66.84 81.29 55.41 62.95
SESS 71.37 63.35 66.34 87.17 67.45 81.75 56.08 62.74
LESS 71.15 64.36 65.33 86.85 67.32 81.61 54.83 62.51
SEC 69.48 63.82 67.27 87.63 66.90 82.03 55.69 63.10
SSEC 70.07 65.21 67.72 87.28 67.02 81.37 56.42 63.25
IPMI 73.12 65.91 68.89 88.01 68.25 82.78 56.47 63.91
PMI (Ours) 91.01 75.22 73.10 89.08 69.03 83.20 57.78 64.25
WCKmeans None 89.80 72.38 65.66 87.84 64.85 82.17 57.22 64.12
ESS 90.11 73.41 64.00 87.52 65.12 81.53 56.30 63.48
SESS 89.98 73.08 63.63 87.41 64.78 81.20 55.98 63.21
LESS 90.11 73.33 64.56 87.00 64.91 81.05 55.62 63.07
SEC 88.85 70.83 63.71 86.90 64.37 80.76 56.35 63.88
SSEC 89.03 70.93 63.63 87.04 64.55 80.92 56.47 63.02
IPMI 90.47 74.03 66.63 87.45 66.13 82.41 57.91 64.33
PMI (Ours) 90.65 74.25 72.25 88.60 66.72 83.07 59.28 64.98
@@ 7.4. Ablation Study and Component Analysis
This section presents ablation studies and the efficacy of each ingredient in Cequel. Due to space constraint, we defer the results of Cequel with various text encoders and LLMs in Appendix D.
Ablation Study. To validate the individual contributions of our key components, we conduct a comprehensive ablation study. Specifically, we explore the effects of the proposed selection algorithms EdgeLLM/TriangleLLM and weighted constrained clustering methods WCKMeans/WCSC. Taking WCSC and TriangleLLM for example, as shown in Table 4, the removal of WCSC results in a decrease in performance, especially on Tweet (ACC falls from 73.10% to 65.33%). This indicates that the weighting of constraints helps in managing noisy or less informative pairs. Additionally, the absence of TriangleLLM also causes a drop in performance, especially on BBC News (ACC declines from 91.01% to 88.72%), highlighting the importance of selecting high-quality constraints. These findings suggest that both components play a significant role.
Edge/Triangle Selection Strategies. We study three strategies in TriangleLLM to select triangles for querying LLMs, where the Max (resp. Min) is to maximize (resp. minimize) the SEC scores, whilst the Random generates triangles randomly. The choice of selection strategy is critical, as it directly determines the quality of constraints obtained within the limited query budget. Table 5 shows that the Max strategy, as implemented by Algorithm 4, achieves the best performance (e.g., 91.01% ACC on BBC News). In comparison, the Min and Random strategies show significantly inferior performance due to selecting low-SEC or random triangles. The results validate the effectiveness of our selection scheme proposed in § 6.
Constrained Clustering Methods. Once the constraints are acquired, a robust constrained clustering algorithm is essential to effectively integrate this guidance into the final partitioning process. Fig. 5 presents the clustering performance of Cequel with various constrained clustering methods, including WCSC (ours), CSC (Wang et al., 2014), PCKMeans (Basu et al., 2004), COPKMeans (Bibi et al., 2023), CCPCKMeans (Rujeerapaiboon et al., 2019), and MPCKMeans (Bilenko et al., 2004) on BBC News, Tweet, Bank77, and Reddit. Our WCSC that assigns personalized weights to constraints achieves the highest accuracy (e.g., 73.10% ACC on Tweet). This demonstrates that WCSC is superior at handling the inherent noise and varying quality of LLM-generated constraints, leading to a more refined and accurate clustering outcome.
Weighting Schemes for Constraints. In Table 6, we report the results of Cequel when using various weighting schemes in our WCSC and WCKmeans on BBC News, Tweet, Bank77, and Reddit. The additional weighting schemes for comparison are listed in the following table:
@@
Table 6 shows that improper weights for constraints lead to a significant adverse impact on the performance in both WCSC and WCKmeans. Amid all weighting schemes, the PMI defined in § 4 consistently achieves the best results. For example, on Tweet, using PMI in WCSC improves over its unweighted version (i.e., None) by a large margin of $7.77\%$, while the best competitor IPMI can only attain an improvement of $3.56\%$.
8. Conclusion
This paper presents Cequel, a framework that cost-effectively leverages LLMs to produce accurate pair-wise constraints for subsequent constrained text clustering. In particular, we develop two greedy algorithms EdgeLLM and TriangleLLM to carefully select text pairs and triplets for querying LLMs so as to obtain the most informative constraints for clustering, and further ameliorate standard constrained clustering algorithms with personalized weights for various pair constraints. Our extensive experiments over benchmark datasets exhibit the high superiority of our proposed techniques in enhancing text clustering quality and reducing the sheer amount of queries needed for LLMs.
Acknowledgements.
This work is partially supported by the National Natural Science Foundation of China (No. 62302414), the Hong Kong RGC ECS grant (No. 22202623), RGC grants 12202024, C1043-24GF, and C2003-23Y, and the Huawei Gift Fund.
GenAI Usage Disclosure
We hereby disclose that GenAI tools were used as a component of the proposed framework in the experimental evaluation.
Appendix A Proof of Lemma 5.2
Proof.
Let $(t_{i},t_{j})=\underset{(t_{a},t_{b})\in\mathcal{S},\ a<b}{\operatorname{arg}\,\operatorname{max}}\;{a}$. In the extreme case, $j=i+1$ and only edges $(t_{a},t_{a+1}),(t_{a},t_{a+2}),\ldots,(t_{a},t_{j})$ are added into $\mathcal{S}$ for each $a\in[1,i-1]$ as they are surely larger than $(t_{i},t_{j})$. Put in another way, for each $t_{a}$ with $a\in[1,i]$, we can select $t_{b}$ with $a+1\leq b\leq j$, i.e., which corresponds to ja=i+1aj-a=i+1-a text instances. Thus, the total number of edges satisfy:
$1+2+\ldots+i=\frac{(i+1)i}{2}=N.$
Then, we can derive
$i^{2}+i-2N=0\Rightarrow i=\frac{\sqrt{8N+1}-1}{2},$
which finishes the proof. ∎
Appendix B Details of Datasets and Baselines
Datasets. The BBC News (Keraghel et al., 2024) dataset contains 2,225 headlines collected from the BBC News website and categorized into 5 high-level topics. It serves as a benchmark for multi-class classification tasks in the news field. The Tweet (Yin and Wang, 2016) dataset consists of 2,472 tweets annotated for relevance to 89 queries from the TREC Microblog track, and is widely used for short-text retrieval and ranking. The Bank77 (Zhang et al., 2021b) dataset includes 3,080 customer service utterances mapped to 77 intent categories, enabling fine-grained intent classification in the financial category. The Reddit (Muennighoff et al., 2023) dataset, with 3,217 posts from online communities, is used primarily for unsupervised topic discovery. The CLINC (Zhang et al., 2022) dataset comprises 4,500 user requests labeled across 10 distinct domains for domain-level text clustering. Finally, the Massive (FitzGerald et al., 2022) dataset contains 11,514 utterances spanning 18 scenario-based categories, and is often used in multilingual and multi-scenario classification settings.
Baselines. We include two sets of baselines in the experiments: embedding-based clustering methods and LLM-assisted clustering methods. In embedding-based clustering methods, we include seven representative baselines that follow a two-stage paradigm: generating fixed-dimensional text embeddings followed by conventional clustering algorithms, such as K-Means++ or spectral clustering (Von Luxburg, 2007). The selected embedding models encompass both traditional and neural models.
  • TF-IDF: A classic sparse representation based on term frequency-inverse document frequency, widely used in earlier text clustering studies.
  • E5 (Wang et al., 2022): A recent retrieval-oriented embedding model pretrained with a multi-task objective, shown to perform well in clustering tasks.
  • DistilBERT and Sentence-BERT (Reimers, 2019): Popular lightweight transformers optimized for sentence-level embeddings.
  • Instructor-Large (Su et al., 2023): A powerful instruction-tuned model designed to align embeddings with various downstream tasks; we use it as the default embedder in most baselines due to its strong performance.
  • OpenAI-GPT (Brown et al., 2020) and LLaMA-2 (7B) (Touvron et al., 2023): Large language models (LLMs) whose embeddings are obtained by averaging final-layer token representations. As LLMs are often optimized for generation, we treat them as black-box embedding extractors here.
These embedding models are coupled with K-Means++ or spectral clustering, depending on the context.
In LLM-assisted clustering methods, we additionally consider three recent strong methods that leverage LLMs beyond embedding extraction.
  • SCCL (Zhang et al., 2021a): A self-supervised contrastive clustering method that improves cluster discrimination by minimizing intra-cluster distance and maximizing inter-cluster margins in embedding space.
  • ClusterLLM (Zhang et al., 2023a): Utilizes LLMs to provide clustering feedback and refines smaller embedding models through LLM guidance. We also include a variant that excludes fine-tuning to assess the impact of feedback alone.
  • PO-PCKMeans (Viswanathan et al., 2023): A semi-supervised approach that constructs pairwise constraints from an LLM based on a small set of labeled examples and integrates them into a constrained clustering framework.
Appendix C Additional Algorithmic Details
C.1. WCKMeans
Input: Corpus $\mathcal{T}$, clusters $K$, MLS $\mathcal{M}$, CLS $\mathcal{C}$, max iterations $T$
Output: Labels $\Gamma$
1
2Normalize $\mathcal{T}$ to get embeddings $\mathbf{X}$;
3 Construct constraint matrix $\mathbf{R}$ with $\mathcal{M}$, $\mathcal{C}$, $|\mathcal{T}|$, and $\mathbf{X}$;
4 Initialize centers $\mathcal{U}\leftarrow\{\mathbf{u}_{1},\ldots,\mathbf{u}_{K}\}$;
5
6while not converged & $t<T$ do
7  foreach $x_{i}\in\mathbf{X}$ do
8
9    $\Gamma_{i}\;\leftarrow\;\underset{1\leq k\leq K}{\operatorname{arg}\,\operatorname{min}}\;\|x_{i}-\mathbf{u}_{k}\|^{2}+\sum\limits_{\begin{subarray}{c}(i,j)\in\mathcal{M}\\ \Gamma_{j}\neq k\end{subarray}}R_{i,j}+\sum\limits_{\begin{subarray}{c}(i,j)\in\mathcal{C}\\ \Gamma_{j}=k\end{subarray}}|R_{i,j}|$
10 Update centers $\mathbf{u}_{k}\leftarrow\texttt{Mean}\bigl{(}\{\,x_{i}\mid\Gamma_{i}=k\}\bigr{)}$ $\forall{k}$;
11
12
return $\Gamma$
Algorithm 5 WCKMeans
Algorithm 5 integrates well-designed weighted constraints into standard K-Means by creating a constraint matrix $\mathbf{R}$, where $R_{i,j}>0$ for must‐links and $R_{i,j}<0$ for cannot‐links. In each iteration, each point $x_{i}$ is assigned to the cluster $k$ that minimizes its squared distance to centroid $\mathbf{u}_{k}$ plus penalties: add $R_{i,j}$ if a must‐link partner $j$ is in a different cluster, and add $|R_{i,j}|$ if a cannot‐link partner $j$ is in the same cluster. After assignments, centroids are updated as the mean of their assigned points. This repeats until convergence or $T$ iterations.
C.2. WCSC
Input: Corpus $\mathcal{T}$, clusters $K$, MLS $\mathcal{M}$, CLS $\mathcal{C}$
Output: Labels $\Gamma$
1
2Normalize $\mathcal{T}$ to get embeddings $\mathbf{X}$;
3 Build affinity matrix $\mathbf{A}$ with $\mathbf{X}$;
4 Construct constraint matrix $\mathbf{R}$ with $\mathcal{M}$, $\mathcal{C}$, $|\mathcal{T}|$, and $\mathbf{X}$;
5 $\mathbf{L}\leftarrow\mathbf{D}^{-1/2}(\mathbf{D}-\mathbf{A})\mathbf{D}^{-1/2},\ \mathbf{R}_{\alpha}\leftarrow\mathbf{D}^{-1/2}\mathbf{R}\mathbf{D}^{-1/2}-\alpha\mathbf{I}$;
6 Compute top-$K$ generalized eigenvectors $\mathbf{V}$ of $\mathbf{L}$ and $\mathbf{R}_{\alpha}$;
7 $\Gamma\leftarrow\texttt{K-Means}(\mathbf{D}^{-1/2}\mathbf{V})$;
8
return $\Gamma$
Algorithm 6 WCSC
Algorithm 6 starts by building a graph that captures how similar each pair of points is in the normalized embedding space. Then selecting top-$K$ eigenvectors finds a low‐dimensional embedding that balances the natural data structure with must‐links and cannot‐links. The $\alpha$ (Wang et al., 2014) serves as a balance factor between the original information captured by the graph and the constraints. Finally, K-Means is applied to produce the cluster labels.
C.3. Edge Query
Input: Corpus $\mathcal{T}$, the number of edges NN, Oracle $\mathbf{O}$
Output: MLS $\mathcal{M}$, CLS $\mathcal{C}$
1
[…]
3 foreach $(t_{i},t_{j})\in\mathcal{S}$ do
4  switch $\mathbf{O}(t_{i},t_{j})$ do
5     case ML do
6       $\mathcal{M}\leftarrow\mathcal{M}\cup\{(t_{i},t_{j})\}$
7    case CL do
8       $\mathcal{C}\leftarrow\mathcal{C}\cup\{(t_{i},t_{j})\}$
9
10
return $\mathcal{M}$, $\mathcal{C}$
Algorithm 7 Edge Query
Algorithm 7 uses Algorithm 1 (GreedyEdgeSelection) to select the top NN instance pairs most informative for querying. Each chosen pair $(t_{i},t_{j})$ is queried to the oracle $\mathbf{O}$, and the corresponding constraints are added based on the responses.
C.4. Triangle Query
Input: Corpus $\mathcal{T}$, the number of triangles $N_{\triangle}$, Oracle $\mathbf{O}$
Output: MLS $\mathcal{M}$, CLS $\mathcal{C}$
1
[…]
3 foreach $(t_{i},t_{j},t_{k})\in\mathcal{S}_{\triangle}$ do
4  switch $\mathbf{O}(t_{i},t_{j},t_{k})$ do
5     case all‐same do
6       $\forall\,(p,q)\in\{(i,j),(i,k),(j,k)\}:\ \mathcal{M}\leftarrow\mathcal{M}\cup\{(t_{p},t_{q})\}$
7    case ij‐same do
8       $\mathcal{M}\leftarrow\mathcal{M}\cup\{(t_{i},t_{j})\},\quad\mathcal{C}\leftarrow\mathcal{C}\cup\{(t_{i},t_{k}),(t_{j},t_{k})\}$
9    case ik‐same do
10       $\mathcal{M}\leftarrow\mathcal{M}\cup\{(t_{i},t_{k})\},\quad\mathcal{C}\leftarrow\mathcal{C}\cup\{(t_{i},t_{j}),(t_{j},t_{k})\}$
11    case jk‐same do
12       $\mathcal{M}\leftarrow\mathcal{M}\cup\{(t_{j},t_{k})\},\quad\mathcal{C}\leftarrow\mathcal{C}\cup\{(t_{i},t_{j}),(t_{i},t_{k})\}$
13    case all‐diff do
14       $\forall\,(p,q)\in\{(i,j),(i,k),(j,k)\}:\ \mathcal{C}\leftarrow\mathcal{C}\cup\{(t_{p},t_{q})\}$
15
16
return $\mathcal{M}$, $\mathcal{C}$
Algorithm 8 Triangle Query
Algorithm 8 utilizes Algorithm 4 (GreedyTriangleSelection) to select $N_{\triangle}$ triangles $(t_{i},t_{j},t_{k})$. For each triangle, the oracle $\mathbf{O}$ returns five outcomes—“all‐same”, “ij‐same”, “ik‐same”, “jk-same”, or “all‐diff”, and the corresponding must‐link or cannot‐link constraints are added to $\mathcal{M}$ and $\mathcal{C}$.
Table 7. Clustering performance with various text encoders
Encoder Method BBC News Tweet Bank77 Reddit
ACC NMI ACC NMI ACC NMI ACC NMI
Sentence-BERT SCCL 82.00 58.33 53.60 72.80 27.80 53.21 36.60 45.71
ClusterLLM 80.08 54.13 57.47 83.74 58.61 77.06 36.09 44.74
PO-PCKMeans 82.70 57.43 61.39 86.12 57.73 78.09 38.48 48.56
Cequel 84.45 61.65 78.48 89.30 67.05 80.46 40.94 49.32
Instructor-Large SCCL 83.60 62.00 36.50 68.50 36.00 61.30 28.70 36.00
ClusterLLM 88.40 69.73 65.81 88.61 68.27 82.45 54.80 62.47
PO-PCKMeans 89.53 72.16 65.58 88.20 67.79 81.92 55.10 62.51
Cequel 91.01 75.22 73.10 89.08 69.03 83.20 57.78 64.25
Appendix D Additional Experiments
Various Text Encoders. Table 7 reports the clustering performance of Cequel and strong baselines using two representative encoders: Sentence-BERT and Instructor-Large. Results demonstrate the consistent superiority of Cequel across datasets and embeddings. For instance, with Instructor-Large on Tweet, Cequel reaches 73.10% ACC and 89.08% NMI, exceeding ClusterLLM by 7.29% and 0.47%. With Sentence-BERT on BBC News, Cequel achieves 84.45% ACC and 61.65% NMI, outperforming PO-PCKMeans by 1.75% and 4.22%, respectively. These gains are consistent across diverse corpora, from structured news to informal tweets and task-oriented texts like Bank77 and Reddit. The results highlight the adaptability of Cequel and its ability to take advantage of powerful pre-trained encoders and informative constraints without task-specific tuning.
$GPT-4o$-Mini